Chapter 9, Section 1, Exercise 002 Use the computer output to estimate the intercept β0 and the slope β- The regression equation is Y 823 - 3.67X. Predictor Coef SE Coef Constant 822.899 89.45 9.20 0.000 -3.674 1.205-3.05 0.006 intercept β0 : Slope 1Chapter 9, Section 1, Exercise 006 tem ute autput a rting a simpla inir medal gi en elo Th2 regressor cquation Y = 81.8-0.0151x. State :he walua af the Sandle ฝ*pe ar the gi en madal n testing if t e il pe in t Fopulation is eren: em zera ide tify the -va and use it and a gni anc숱 eve tú make a dea case aer about the a ective ess oft/瑜广edel. Constant 84.38 12.16 6.97 0.000 -0.01539 001283-1.20 .245 Sample skpe: galle Is the modl Mective? N ou would like to Shaw Work for thisquestion onSw wlark
Answer #1

Q.1) The regression equation is,

Y = 823 - 3.67X

Where,

Intercept eta_0 : 823

Slope 61 : -3.67

Q.2) The regression equation is,

Y = 93.3 - 7.89X

Sample slope = -7.89

p-value = 0.000

Here, p-value = 0.000 < a-0.05

Implies that model is effective.

Q.3) The regression equation is,

Y = 84.8 - 0.0154X

Sample slope = -0.0154

p-value = 0.245

Here, p-value = 0.245 > a-0.05

Implies that model is not effective.